OpenStudy (anonymous):

MatLab Is this algorithm correct for solving the heat equation with a Crank-Nicolson scheme? Thanks!

OpenStudy (anonymous):

clc; clear; n = 100; m = 100; dt = 0.0000049; dx = 0.01; k = 1; s = k * dt / ( dx ^2 ); ui = 10; ud = 5; A = zeros( n, m ); A( 1, 1 ) = ( 1 + s ); A( 1, 2 ) = -s / 2; A( n, m ) = ( 1 + s ); A( n, m - 1 ) = -s / 2; for i = 2 : n - 1 A( i, i - 1 ) = -s / 2; A( i, i ) = ( 1 + s ); A( i, i + 1 ) = -s / 2; end u = zeros( n, 1 ); d = zeros( n, 1 ); d( 1 ) = ui; d( n ) = ud; for i = 2 : n - 1 d( i ) = s / 2 * u ( i + 1 ) + ( 1 - s ) * u( i ) + s / 2 * u ( i - 1 ); end for i = 1 : m u = A \ d; d = u; d( 1 ) = ui; d( n ) = ud; end